OpenStudy (asylum15):

Integration substitution:

OpenStudy (asylum15):

integral of \[\frac{ -x^4 }{ \sqrt{x^5+2} } dx\]

OpenStudy (asylum15):

Can someone explain step by step how this substitution works please? Thank you.

OpenStudy (anonymous):

assume x^5+2 =z^2 so that 5x^4 dx = 2z dz and then proceed further

OpenStudy (asylum15):

Let u = x^5 + 2 du/dx = 5x^4 du = 5x^4 . dx 1/5 du = x^4 . dx

OpenStudy (asylum15):

How does this substitute BACK into the original equation? This is what I don't understand.

OpenStudy (anonymous):

sincethere is a sqrt in the denominator it would be better to assume the denominator as some perfect squares

OpenStudy (asylum15):

Any input Zak?

OpenStudy (asylum15):

Oh NO! I clicked close accidently :(

OpenStudy (anonymous):

assume x^5+2 =z^2 so that 5x^4 dx = 2z dz so the required integration can be re written as integration of -(1/5) *(5x^4) dx/(sqrt(x^5+2) on substitution x^5+2 =z^2 and 5x^4 dx = 2z dz we have the question as (-1/5) 2zdz/sqrt(z^2) thus we e now the integration of (-2/5) dz which on integration gives (-2/5) z + c now we have z= sqrt(x^5+2) hence our answer is (-2sqrt(x^5+2)/5) + c

OpenStudy (asylum15):

Thanks! :)

OpenStudy (anonymous):

welcome dear well i hope u have understood